OpenStudy (anonymous):

solve this by using the elimination method 2x + 5y = 13 3x - 4y = 8

HanAkoSolo (jamierox4ev3r):

multiply all of the values in the upper equation by 3, and multiply all the values in the lower equation by -2.

HanAkoSolo (jamierox4ev3r):

Then the two equations become: 6x+15y=39 -6x+8y=-16

HanAkoSolo (jamierox4ev3r):

Based from this, can you solve?

HanAkoSolo (jamierox4ev3r):

@KateNicole ?

OpenStudy (anonymous):

sorry, no probably not, I need to see all of the work to understand, and since the class is online I usually dont get most of it at first

HanAkoSolo (jamierox4ev3r):

ah i see okay

HanAkoSolo (jamierox4ev3r):

then, i will show you the rest of the process: 6x+15y=39---------> 23y=23 -6x+8y=-16 --------> y=1

HanAkoSolo (jamierox4ev3r):

now that we know that y=1, we can use this to find the x-variable

HanAkoSolo (jamierox4ev3r):

2x+5(1)=13 2x+5=13 -5 -5 2x=8 /2 /2 x=4 3x-4(1)=8 3x-4=8 +4 +4 3x=12 /3 /3 x=4

HanAkoSolo (jamierox4ev3r):

by proving both of the equations, we have confirmed that y=1 and x=4

HanAkoSolo (jamierox4ev3r):

Does that make sense? @KateNicole ?

OpenStudy (anonymous):

More sense then it made before, thanks, I can probably do the rest of them now:)

HanAkoSolo (jamierox4ev3r):

im glad :)